### v1

@@ 3.2 Random Forests
Random Forests (RFs) (Breiman, 2001) learn relationships between feature vectors and response variables. Training data comprising a collection of feature vectors $X_{i}\in\mathcal{X}$ and corresponding response variables $y_{i}\in\mathcal{Y}$ are used to train a RF, enabling it to predict the true response variable $y^{*}$ of an unseen feature vector $X^{*}$. Regression RFs are used when $y_{i}$ are continuous values and classification RFs are used when $y_{i}$ are discrete labels. Raynal et al. (2018) used a regression RF for parameter inference by using simulated data $\mathcal{D}_{i}$ to learn the relationship between summary statistics $X_{i}$ and parameter values $y_{i}=\theta_{i}$. Given summary statistics $X^{*}$ of unseen data $\mathcal{D}^{*}$, the prediction $RF(X^{*})$ predicts the true parameter value $\theta^{*}$. In addition to predicting unseen feature vectors, a trained RF provides useful information about the training data. The out-of-bag prediction $RF_{\text{oob}}(X_{i})$ estimates the (known) response variable $y_{i}$ using pairs from the training data other than $(X_{i},y_{i})$. The out-of-bag error rate $p(RF(X_{i})\neq y_{i})$ then gives a (unbiased) measure of how well the relationship between $X_{i}$ and $y_{i}$ is captured by the rest of the training data. A trained RF also gives a measure of the importance of each co-variate $j$ within feature vectors $X_{i}=(X_{i}^{0},\dots,X_{i}^{j},\dots,X_{i}^{n_{\text{f}}})$ to the problem of predicting response variable $y_{i}$. Intuitively, important features are those whose values within $X_{i}$ and $X_{i}^{\prime}$ differ when $y_{i}$ and $y_{i}^{\prime}$ do, and which are hence useful in learning the relationship between training data $\mathcal{X}$ and $\mathcal{Y}$.
@@ 3.3 Model Selection
Given observed data $\mathcal{D}^{*}$, the model posterior $p(m|\mathcal{D}^{*})$ is the probability that models $m=m_{i}$ generated $\mathcal{D}^{*}$ for $m_{i}=1,2,\dots$. Algorithm 1 and its extensions rely on the approximation $p(\Theta|X^{*})\approx p(\Theta|\mathcal{D}^{*})$, which holds as long as the vector $X_{i}$ carries a similar amount of information about the parameter value $\theta_{i}$ as the simulated data $\mathcal{D}_{i}$ itself. However, the information loss suffered by a collection of summary statistics may vary between models (Robert et al., 2011), so it is inadvisable simply to infer $m_{i}$ as a (discrete) parameter using Algorithm 1. Pudlo et al. (2015) instead used two RFs to approximate $p(m|\mathcal{D}^{*})$. A classification RF learns the relationship between simulated data $X_{i}$ and model label $y_{i}=m_{i}$ and gives a prediction $RF(X^{*})$ of the model $m^{*}$ which generated unseen data $\mathcal{D}^{*}$. A regression RF is then trained to learn the relationship between $X_{i}$ and $p(RF_{\text{oob}}(X_{i})\neq m_{i})$–the out-of-bag error rate of the classification RF. The regression RF is then used to estimate posterior probability $p(m=m^{*}|\mathcal{D}^{*})$ as $1-p(RF(X^{*})\neq m^{*})$.
@@ 4 Spatial Parameter Inference and Model Selection
Given observed data $\mathcal{D}^{*}$, we wish to approximate the parameter posterior $p(\Theta|\mathcal{D}^{*})$ for candidate models $m=m_{i}$ for $i=1,2,\dots$ and the model posterior $p(m|\mathcal{D^{*}})$. Informative summary statistics may be used to define a distance function $\nu(\mathcal{D}^{*},\mathcal{D}_{i})$ to infer parameters using ABC. However, uninformative or poorly scaled summary statistics may misrepresent the difference between datasets generated by similar parameters (Blum et al., 2013). We therefore seek a collection of summary statistics that quantify simulated data and, in particular, quantify how simulated data changes when different model rules and parameters are used to generate it. We use informative summary statistics to approximate parameter and model posteriors in a three-step pipeline. We test this pipeline on toy models in Supplementary Information Section 10 and apply it to the three angiogenesis models in Section 5.
Step 1: Identify informative summary statistics for each model
We draw parameter values $\theta_{i}$ from the prior distribution $p(\Theta)$ for each parameter in each model, simulate model data $\mathcal{D}_{i}$, and compute spatially-averaged and topological statistics $X_{i}$ from the final simulated time-step. We train regression RFs to learn the relationship between summary statistics $X_{i}$ and parameter values $\theta_{i}$–one RF for each parameter in each model. We then rank the spatially-averaged and topological statistics by their importance according to the RF (see Section 3.2). In each RF, feature importance decreases exponentially (as in Raynal et al. (2018)) and a small subset of summary statistics provides most of the predictive power of each RF. We select an equal number of informative summary statistics from each RF and collect a total of $n_{\text{s}}$ for each model. We find that $n_{\text{s}}=100$ is a good choice for the AC, SL and PS angiogenesis models. See Supplementary Information Section 9 for a full definition of RF feature importance and a discussion of how we choose $n_{\text{s}}$.
@@
We use the distance function $\nu(\mathcal{D}^{*},\mathcal{D}_{i})=\norm{x^{*}-x_{i}}_{2}$ to compare observed data $\mathcal{D}^{*}$ to simulated data $\mathcal{D}_{i}$. Here, $x_{i}$ is the vector of summary statistics $X_{i}$ restricted to the top $n_{\text{s}}$ features for each model, $x^{*}$ is computed from observed data, and the distance is averaged over multiple instances of observed data. We scale each summary statistic by the largest absolute value of that summary statistic in the training data. Doing so ensures that each summary statistic contributes approximately equally to the distance function $\nu$ and limits the influence of poorly scaled summary statistics. We then use the ABC-SMC algorithm of Del Moral et al. (2012) to approximate $p(\Theta|\mathcal{D}^{*})$ for each model.
@@
Following Pudlo et al. (2015), we train a classification RF to learn the relationship between (unscaled) summary statistics $X_{i}$ and model indices $m_{i}$ in the training data. We modify $X_{i}$ to contain only those summary statistics which appear among the $n_{\text{s}}$ most important summary statistics for all models. We then train a regression RF to learn the relationship between $X_{i}$ and $p(RF_{\text{oob}}(X_{i})\neq m_{i})$–the probability that the predicted model index is incorrect. The classification RF gives an estimate $RF(X^{*})$ of the model $m^{*}$ that generated the observed data $\mathcal{D}^{*}$, and the regression RF is used to estimate $p(m=m^{*}|\mathcal{D}^{*})$.
Figure 3: We approximate the model posterior $p(m|\mathcal{D}^{*})$ using the same four test-cases from Figure 2 (again highlighting one example of observed data $\mathcal{D}^{*}$). For each model, we show one example of simulated data associated to an inferred parameter in that model’s approximate parameter posterior. Each ‘prediction’ shows an example of that model’s approximation of the true data generation process.
@@ 5.1 RFs find small subsets of informative summary statistics
We sample $n=10,000$ model parameters from uniform priors with ranges taken from existing literature, or by analysing each model’s data generation rules (See Supplementary Information Section 7 for details). We select $n_{\text{s}}=100$ summary statistics for each model and report the type of summary statistics selected–spatially-averaged or topological, computed in the horizontal or vertical direction–in Figure 1. A mixture of spatially-averaged and topological statistics are selected for each parameter, however there is a clear preference for topological statistics. More summary statistics computed in the vertical direction are selected than those computed in the horizontal direction, which is unsurprising, since most parameters moderate the movement of ECs upwards. Two exceptions are the randomness parameter $\sigma$ of the SL model and the turning coefficient $D_{r}$ in the PS model. These parameters cause ECs to deviate from their upward trajectory, and their RFs rely on summary statistics computed in the horizontal direction for inference.
@@ 5.2 ABC-SMC infers four parameters for each model and reproduces observed data
We create two test-cases for each model by simulating data at known parameter values $10$ times. Each test-case exhibits quantitatively different vascular networks that each model can produce. We use the $100$ summary statistics reported in Figure 1 to define a distance function between simulated and observed data as described in Section 3.1, averaging this distance across the $10$ instances of observed data. We use ABC-SMC to infer four parameters for each model and show the resulting approximate posteriors in Figure 2. In each test-case, the approximate posterior is unimodal and encompasses the true parameter, often close to its densest part.
@@ 5.3 Random Forests correctly select models
Using the same test-cases, we perform Step 3, training a classification RF to predict the model index of unseen test data $\mathcal{D}^{*}$ and a regression RF to predict the probability of mis-classifying each test-case. Figure 3 shows the resulting approximate model posteriors, which identify the correct model with high probability in each test-case. We simulate each model at a parameter value drawn from its approximate parameter posterior and show the resulting ‘prediction’, which is a sample from the model’s approximation of the true data generating process. Each model generates data that is visually similar to the observed data, however we can identify the true model in each test-case.
@@ 6 Discussion
Using Topological Data Analysis (TDA) and Approximate Bayesian Computation (ABC), we have developed a pipeline for parameter inference and model selection applicable to complex spatial models. In previous work, TDA characterised the effect of two parameters in the AC model (Nardini et al., 2021) and was used along with ABC to infer them (Thorne et al., 2022). We extend this work by identifying a subset of informative summary statistics from multiple topological filtrations and using them to infer four parameters in three angiogenesis models using ABC-SMC. We further show how RFs can be used with TDA to approximate model posteriors and compare candidate models.
While we validated our pipeline using synthetic data simulated from angiogenesis models, several previous studies used in vitro data to inform model rules and parameters (Milde et al., 2013; Connor et al., 2015; Vergroesen et al., 2024). ABC provides a robust statistical framework for learning parameters and evaluating models, and TDA provides a variety of filtrations and vectorisations which may be adapted to different spatial data. The present work would therefore enhance previous model analysis, and in future work we will apply our methods to experimental data.
In this study, we considered three models in which the paths traced by tip EC form a static blood vessel network. In reality, sustained proliferation and vessel remodelling, where vasculature continually evolves after it is laid down, is characteristic of tumour-induced angiogenesis (Farnsworth et al., 2014). Indeed, vascular renormalisation, in which vessel-targetting agents prune small or inefficient blood vessels, is a theorised treatment strategy (Magnussen and Mills, 2021) aiming to temporarily enhance perfusion of the tumour to increase the effectiveness of radiotherapy (Köry et al., 2024). Topological invariants taking into account time-evolving data (Carlsson and de Silva, 2010), directed flow networks (Chaplin et al., 2024) and multiparameter filtrations (Vipond, 2020) could quantify such structural changes over time and be used to calibrate more sophisticated models.
Although we specialised our pipeline to discrete models of tumour-induced angiogenesis, its flexibility allows application to a range of spatio-temporal models. Any summary statistic which aims to quantify the desired properties of simulated data would be identified by the RF in Step 1 if it captures the effect of changing model parameters. In future, we will use our pipeline to robustly compare continuum, cell-based, agent-based, and discrete models by their ability to reproduce observed data.
@@ 7 Angiogenesis Models
We use the Anderson-Chaplain (AC) model (Anderson and Chaplain, 1998), the Stokes-Lauffenberger (SL) model (Stokes et al., 1991), and the Plank-Sleeman (PS) model (Plank and Sleeman, 2004) as our case-study for topological model selection. In this section, we present the models in dimensionless form, giving formulae and simulation details for each model. We list model parameters in Table 1 and provide illustrative schematics in Figure 4.
Each model simulates the movement of multiple tip endothelial cells (ECs) in a two-dimensional square domain =$\mathcal{I}=\{(x,y):x,y\in[0,1]\}\subset\mathbb{R}^{2}$. The position of a tip EC at time $t$ is described by a variable $s^{t}=(s_{x}^{t},s_{y}^{t})\in\mathcal{I}$. We initialise four tip ECs along the lower edge of the domain (furthest from the tumour). We create four variables $s$ such that ${s}^{0}=(-1/8+\upsilon/4,0)$ for $\upsilon=1,2,3,4$. At discrete time intervals $\Delta t$, tip ECs migrate to a new position $s^{t+\Delta t}\in\mathcal{I}$, which is determined by a set of model-specific movement rules. In each model, we assume that a tumour is located along the domain’s upper boundary ($y=1$), acting as a source of vascular endothelial growth factors (VEGF). For simplicity, we prescribe the initial profile of VEGF $c(x,y,t=0)=y$ in which the VEGF concentration decreases linearly with distance from the tumour. When a tip EC moves from $s^{t}$ and $s^{t+\Delta t}$, the line segment connecting subsequent positions is assumed to be occupied thereafter by immobile stalk EC (the snail-trail model (Balding and McElwain, 1985)). Each model simulates EC movement over the (dimensionless) time interval $\mathcal{T}=[0,t_{\text{final}}]$. We set $\Delta t$ = 0.01 and $t_{\text{final}}=4$ meaning that model simulations terminate after $400$ time-steps.

### v2

@@ 3.2 Random Forests
Random Forests (RFs) (orginalRF) learn relationships between feature vectors and response variables. Training data comprising a collection of feature vectors $X_{i}\in\mathcal{X}$ and corresponding response variables $y_{i}\in\mathcal{Y}$ are used to train a RF, enabling it to predict the true response variable $y^{*}$ of an unseen feature vector $X^{*}$. Regression RFs are used when $y_{i}$ are continuous values and classification RFs are used when $y_{i}$ are discrete labels. ABCRF_paraminf used a regression RF for parameter inference by using simulated data $\mathcal{D}_{i}$ to learn the relationship between summary statistics $X_{i}$ and parameter values $y_{i}=\theta_{i}$. Given summary statistics $X^{*}$ of unseen data $\mathcal{D}^{*}$, the prediction $RF(X^{*})$ predicts the true parameter value $\theta^{*}$. In addition to predicting unseen feature vectors, a trained RF provides useful information about the training data. The out-of-bag prediction $RF_{\text{oob}}(X_{i})$ estimates the (known) response variable $y_{i}$ using pairs from the training data other than $(X_{i},y_{i})$. The out-of-bag error rate $p(RF(X_{i})\neq y_{i})$ then gives a (unbiased) measure of how well the relationship between $X_{i}$ and $y_{i}$ is captured by the rest of the training data. A trained RF also gives a measure of the importance of each co-variate $j$ within feature vectors $X_{i}=(X_{i}^{0},\dots,X_{i}^{j},\dots,X_{i}^{n_{\text{f}}})$ to the problem of predicting response variable $y_{i}$. Intuitively, important features are those whose values within $X_{i}$ and $X_{i}^{\prime}$ differ when $y_{i}$ and $y_{i}^{\prime}$ do, and which are hence useful in learning the relationship between training data $\mathcal{X}$ and $\mathcal{Y}$.
@@ 3.3 Model Selection
Given observed data $\mathcal{D}^{*}$, the model posterior $p(m|\mathcal{D}^{*})$ gives the probability that models $m=m_{i}$ generated $\mathcal{D}^{*}$. ABC algorithms rely on the approximation $p(\Theta|X^{*})\approx p(\Theta|\mathcal{D}^{*})$, which holds as long as the vector $X_{i}$ carries a similar amount of information about the parameter value $\theta_{i}$ as the simulated data $\mathcal{D}_{i}$ itself. However, the information loss suffered by a collection of summary statistics may vary between models (noConfidenceABCModelChoice), so it is inadvisable simply to infer $m_{i}$ as a (discrete) parameter using an ABC algorithm. ABCRF_modelchoice instead used two RFs to approximate $p(m|\mathcal{D}^{*})$. A classification RF learns the relationship between simulated data $X_{i}$ and model label $y_{i}=m_{i}$ and gives a prediction $RF(X^{*})$ of the model $m^{*}$ which generated unseen data $\mathcal{D}^{*}$. A regression RF is then trained to learn the relationship between $X_{i}$ and $p(RF_{\text{oob}}(X_{i})\neq m_{i})$–the out-of-bag error rate of the classification RF. The regression RF is then used to estimate posterior probability $p(m=m^{*}|\mathcal{D}^{*})$ as $1-p(RF(X^{*})\neq m^{*})$.
@@ 4 Spatial Parameter Inference and Model Selection
Given observed data $\mathcal{D}^{*}$, we wish to approximate the parameter posterior $p(\Theta|\mathcal{D}^{*})$ for candidate models $m=m_{1},m_{2},\dots$ and the model posterior $p(m|\mathcal{D^{*}})$. Informative summary statistics may be used to infer parameter values using ABC, but uninformative or poorly scaled summary statistics may misrepresent the difference between datasets generated by similar parameters (summaryStatisticChoice). We therefore seek a collection of summary statistics that quantify simulated data and, in particular, quantify how simulated data changes when different model rules and parameters are used to generate it. We use informative summary statistics to approximate parameter and model posteriors in a three-step pipeline. We test this pipeline on toy models in Supplementary Information section 10 and apply it to the three angiogenesis models in section 5.
Step 1: Identify informative summary statistics
We use RFs to find a small subset of summary statistics to be used in ABC. To generate training data, we draw parameter values $\theta_{i}$ from the prior distribution $p(\Theta)$ for each parameter in each model, simulate model data $\mathcal{D}_{i}$, and compute spatially-averaged and topological summary statistics $X_{i}$ from the final simulated time-step. We train regression RFs to learn the relationship between summary statistics $X_{i}$ and parameter values $\theta_{i}$–one RF for each parameter in each model. We then rank the spatially-averaged and topological summary statistics by their importance according to the RF (see section 3.2). In each RF, feature importance decreases exponentially (as in ABCRF_paraminf) and a small subset of summary statistics provides most of the predictive power of each RF. We select an equal number of informative summary statistics from each RF, collecting a total of $n_{\text{s}}=100$ for each model. See Supplementary Information section 9 for a full definition of RF feature importance and a discussion of how we choose $n_{\text{s}}$. RFs identify which summary statistics quantify the effect of each parameter on simulated data and allow us to omit those summary statistics which do not.
@@
We use the summary statistics identified by step 1 to define a distance function for use in ABC. We use $\nu(\mathcal{D}^{*},\mathcal{D}_{i})=\norm{x^{*}-x_{i}}_{2}$, where $x_{i}$ is the vector $X_{i}$ restricted to the top $n_{\text{s}}$ summary statistics identified in step 1, $x^{*}$ is computed from observed data, and the distance is averaged over multiple instances of observed data. We scale each summary statistic by the largest absolute value of that summary statistic in the training data. We then use the ABC-SMC algorithm of smcDelMoral to approximate $p(\Theta|\mathcal{D}^{*})$ for each model. By using only those summary statistics which quantify the effect of parameter values on simulated data, we ensure $\nu$ is informative about the value of $\theta$ used to generate the observed data. Scaling ensures that each summary statistic contributes approximately equally to the distance function $\nu$ and limits the influence of poorly scaled summary statistics.
@@
Using summary statistics which are informative for all models, we use two more RFs to estimate the model posterior. Following ABCRF_modelchoice, we train a classification RF to learn the relationship between (unscaled) summary statistics $X_{i}$ and model indices $m_{i}$ in the training data. We modify $X_{i}$ to contain only those summary statistics which appear among the $n_{\text{s}}$ most important summary statistics for all models. We then train a regression RF to learn the relationship between $X_{i}$ and $p(RF_{\text{oob}}(X_{i})\neq m_{i})$–the probability that the predicted model index is incorrect. The classification RF gives an estimate $RF(X^{*})$ of the model $m^{*}$ that generated the observed data $\mathcal{D}^{*}$, and the regression RF is used to estimate $p(m=m^{*}|\mathcal{D}^{*})$ as $1-p(RF_{\text{oob}}(X^{*})\neq m^{*})$. We choose the value of $n_{\text{s}}$ in step 2 large enough to ensure that some informative summary statistics are selected for all three models under consideration–we use these to approximate the model posterior.
@@ 5.1 RFs find small subsets of informative summary statistics
We sample $n=10,000$ model parameters from uniform priors with ranges taken from existing literature, or by analysing each model’s data generation rules (See Supplementary Information section 7 for details). We select $n_{\text{s}}=100$ summary statistics for each model and report the type of summary statistics selected–spatially-averaged or topological, computed in the horizontal or vertical direction–in Figure 1. A mixture of spatially-averaged and topological summary statistics are selected for each parameter, however there is a clear preference for topological summary statistics for each parameter. In general, summary statistics computed in the vertical direction are selected more often than those computed in the horizontal direction, which is unsurprising, since most parameters regulate the movement of ECs upwards. However, the randomness parameter $\sigma$ of the SL model and the turning coefficient $D_{r}$ of the PS model are two exceptions. These parameters cause ECs to deviate from their upward trajectory, and their inference therefore relies on topological summary statistics computed in the horizontal direction. No spatially-averaged summary statistics computed in the horizontal ($x$) direction are chosen, indicating that these measures are too coarse to distinguish different model simulations.
Figure 4: We approximate the model posterior $p(m|\mathcal{D}^{*})$ using the six test-cases from Figure 3, highlighting one example of $\mathcal{D}^{*}$. For each test-case, we show one example of data simulated using an inferred parameter from each model’s approximate parameter posterior. Each ‘prediction’ shows an example of that model’s approximation of the true data generation process.
@@ 5.2 ABC-SMC infers four parameters for each model and reproduces observed data
We create two test-cases for each model by simulating data at known parameter values $10$ times. Each test-case exhibits quantitatively different vascular networks that each model can produce. Following steps 1 and 2 of section 4, we use ABC-SMC to infer four parameters for each model and show the resulting approximate posteriors in Figure 3. In each test-case, the approximate posterior is unimodal and encompasses the true parameter, often close to its densest part.
@@ 5.3 Random Forests correctly select models
Using step 3 of section 4, we approximate the model posterior for each of the six test-cases. Figure 4 shows the resulting approximate model posteriors, which identify the correct model with high probability in each test-case. We simulate each model at a parameter value drawn from its approximate parameter posterior and show the resulting ‘prediction’, which is a sample from the model’s approximation of the true data generating process. Each model generates data that is visually similar to the observed data, however we can identify the true model in each test-case.
@@ 6 Discussion
Using Topological Data Analysis (TDA) and Approximate Bayesian Computation (ABC), we have developed a pipeline for parameter inference and model selection applicable to complex spatial models. In previous work, TDA characterised the effect of two parameters in the AC model (nardiniclustering) and was combined with ABC to infer them (tabc). Here we extend this work by identifying a subset of informative summary statistics from multiple topological filtrations and use them to infer four parameters in three angiogenesis models using ABC-SMC. We further show how RFs can be used with TDA to approximate model posteriors and compare candidate models.
While we validated our pipeline using synthetic data simulated from angiogenesis models, previous studies used in vitro data to inform model rules and parameters (mildeAngio; connorAngio; merksAngioModels). The present work therefore enhances previous model analysis, since ABC provides a statistical framework for learning parameters and evaluating models, and TDA provides a variety of filtrations and vectorisations which may be adapted to different spatial data. In future work we will apply our pipeline to real experimental data, which will take the place of the synthetic test-cases in Figures 3 and 4 wherein the true parameter values and correct model will not be known.
In this study, we considered three models in which the paths traced by tip ECs form a static blood vessel network. Topological features, therefore, evolve monotonically in time in these models, and computing EPH at the final timestep was sufficient to infer parameters. In reality, sustained proliferation and vessel remodelling, where vasculature continually evolves after it is laid down, is characteristic of tumour-induced angiogenesis (vascularremodeling). Indeed, vascular renormalisation, in which vessel-targetting agents prune small or inefficient blood vessels, is a theorised treatment strategy (vascularnormalisation) aiming to temporarily enhance perfusion of the tumour to increase the effectiveness of radiotherapy (jain_antiangio). Topological invariants that account for time-evolving data (crokervineyardmulti), directed flow networks (grpph) and multiparameter filtrations (mpp) could quantify such structural changes over time and be used to calibrate more sophisticated models. Rather than fixing the duration of simulations in these cases, simulation time could be an extra parameter to be inferred by our pipeline.
Although we specialised our pipeline to discrete models of tumour-induced angiogenesis, its flexibility allows application to a range of spatio-temporal models. Any summary statistic which quantifies the desired properties of simulated data would be identified by the RF in step 1 if it captures the effect of changing model parameters. In future, we will use our pipeline to systematically compare continuum, cell-based, agent-based, and discrete models by their ability to reproduce observed data.
@@ 7 Angiogenesis Models
We use the Anderson-Chaplain (AC) model (andersonChaplain), the Stokes-Lauffenberger (SL) model (stokesLauffenberger), and the Plank-Sleeman (PS) model (plank_sleeman) as our case-study for topological model selection. In this section, we present the models in dimensionless form, giving formulae and simulation details for each model. We list model parameters in Table 1 and provide illustrative schematics in Figure 5.
Each model simulates the movement of multiple tip endothelial cells (ECs) in a two-dimensional square domain $\mathcal{I}=\{(x,y):x,y\in[0,1]\}\subset\mathbb{R}^{2}$. The position of a tip EC at time $t$ is described by a variable $s^{t}=(s_{x}^{t},s_{y}^{t})\in\mathcal{I}$. We initialise four tip ECs along the lower edge of the domain (furthest from the tumour). We create four variables $s$ such that ${s}^{0}=(-1/8+\upsilon/4,0)$ for $\upsilon=1,2,3,4$. At discrete time intervals $\Delta t$, tip ECs migrate to a new position $s^{t+\Delta t}\in\mathcal{I}$, which is determined by a set of model-specific movement rules. In each model, we assume that a tumour is located along the domain’s upper boundary ($y=1$), acting as a source of vascular endothelial growth factors (VEGF). For simplicity, we prescribe the initial profile of VEGF $c(x,y,t=0)=y$ in which the VEGF concentration decreases linearly with distance from the tumour. When a tip EC moves from $s^{t}$ and $s^{t+\Delta t}$, the line segment connecting subsequent positions is assumed to be occupied thereafter by immobile stalk EC (the snail-trail model (originalSnailTrail)).